# CVBench: Evaluating Cross-Video Synergies for Complex Multimodal Understanding and Reasoning

### v2

@@ 4.2 Results and Findings
Closed-source vs. open-source models.
Closed-source models consistently outperform open-source models. This suggests that proprietary architectures may utilize more refined optimization techniques or superior capabilities in handling complex cross-video relationships, revealing a substantial gap between commercial-grade and publicly available MLLMs in this domain.
Run ID #1 #2 #3 Avg.
@@
Table 5: Ablation on input perturbations.
Sensitivity to input perturbations.
@@
Table 6: Impact of frame number & resolution.
Frame resolution.
We investigate the impact of frame resolution using Qwen2.5-VL, considering different resolutions and varying numbers of input frames. With 8 frames, accuracy increases slightly from 46.0% at 224×224 to 46.8% at 448×448 (+0.8%), indicating that higher resolution leads to better performance. Similar trends are observed with 16 and 32 frames. Additionally, increasing the number of input frames consistently improves performance across both resolutions.
Figure 4: Impact of video number and duration.
Number of video instances & duration.
Fig. 4 illustrates how video duration and the number of videos affect model performance (Qwen2.5-VL). Accuracy improves with video length up to 900 seconds (48.8% for 0–300s), but declines for longer videos (45.6% for 900–1200s), likely due to increased sequence complexity and context management challenges. We also highlight the difficulty of multi-video reasoning. Models that perform well on single-video tasks struggle with cross-video tasks. Shown in Fig. 4, the model achieves 52.2%/47.9%/44.8% accuracy with two/three/four videos, showing increased complexity of multi-video tasks.
Qualitative analysis of reasoning-based MLLMs.
Fig. 5 illustrates the reasoning process of Video-R1, emphasizing the implicit cues that shed light on the model’s behavior. The left sub-figure depicts hallucination in reasoning, where the model generates responses that are logically disconnected. In contrast, the right sub-figure highlights the model’s limitations in achieving deeper understanding and reasoning compared to human cognition. Video-R1 frequently produces logically incorrect responses and struggles to reach human-level reasoning. This shows the need for ongoing advancements in multimodal models to better emulate human-like reasoning, particularly for tasks that require cross-modal integration and contextual understanding in real-world scenarios.
@@ 5 Conclusion
This paper introduces CVBench, a novel benchmark for evaluating the performance of MLLMs in cross-video understanding and reasoning tasks. Our evaluation demonstrates that, while MLLMs excel in single-video tasks, they encounter significant challenges in spatiotemporal alignment, entity recognition, and temporal reasoning when processing multiple video streams. CVBench provides a comprehensive dataset and clearly defined tasks, highlighting the unique demands of cross-video scenarios, which require the synthesis of information across both temporal and spatial dimensions. We present a systematic evaluation of leading MLLMs on CVBench, revealing notable performance gaps and critical challenges in cross-video reasoning, and further identifying key research directions to advance the field. Addressing these challenges is essential for improving multi-video reasoning and expanding the applicability of MLLMs to real-world problems.
Appendix A Dataset Collection
A.1 Task Definitions
@@
While single-video tasks assess a model’s comprehension of individual videos, CVBench evaluates cross-video relational reasoning. This critical capability enables MLLMs to advance from controlled lab settings to robust real-world deployment. CVBench introduces 15 carefully curated evaluation dimensions that collectively measure a model’s capacity for inter-video relational reasoning, each targeting a unique but complementary aspect of cross-video understanding. An overview of these tasks, their concepts, differences, and example questions is provided in Table 7.
Task 1: Cross-video anomaly detection Concept: The comparative analysis of multi-source video data identifies anomalous segments deviating from normal patterns through global spatiotemporal feature and behavioral pattern comparisons. Difference: Analyzes anomaly saliency through multi-video event richness and viewpoint coherence rather than single-video perspectives. Example Questions: Question: Which video features individuals directly engaged in construction work rather than showcasing materials or interior design? Option: A. Video 1, B. Video 2, C. Video 3, D. Video 4 Task 2: Cross-video scene recognition Concept: Fusion of semantic features from multiple videos to locate scene segments matching target spatiotemporal attributes. Difference: Focuses on cross-video differences rather than single-video similarities. Example Questions: Question: “Did both videos feature a team successfully scoring a decisive goal or point during the match’s closing moments?” Option: Yes/No. Task 3: Multi-video key-action recognition Concept: This category involves spatiotemporal alignment of motion trajectories across videos to identify differences in action execution, supporting action quality assessment and optimization. Difference: It emphasizes modeling the spatiotemporal consistency of actions across multiple videos, rather than classifying actions within a single video. Example Questions: Question: What is the key action difference between the presenter in Video 1 and the presenter in Video 2? Option: A. In Video 1, holding products is key, while in Video 2, gesturing with hands dominates., B. In Video 1, gesturing is emphasized more, while in Video 2, interacting with products is dominant., C. In both, sitting is a primary action, but outdoor walking is unique to Video 2., D. Video 2 uses textual prompts for interactions, while Video 1 does not. Task 4: Cross-video procedural transfer Concept: This category focuses on cross-video dynamic knowledge transfer capabilities, this approach analyzes the knowledge in the source video (such as first aid steps or mathematical problem solving methods) and adaptively refines or supplements the execution logic of similar tasks in the target video, while overcoming the interference caused by scene differences. Difference: Associating procedural semantics across videos, as a single video can only capture isolated processes and cannot achieve knowledge generalization or adaptation. Example Questions: Question: What decor themes from Video 2 can be transferred to Video 4 to enhance the bar’s elegance? Option: A. Floral arches and greenery., B. Heart motifs on walls., C. Outdoor nature-inspired pieces., D. Grand tents and fairy lights.
Task 5: Joint-video summarization Concept: This category involves extracting semantic information from multiple videos to generate event logic and descriptions that satisfy specific query requirements. Difference: Cross-video information fusion and semantic compression, rather than simple fragment stitching. Example Questions: Question: How can the overall narrative of Videos 1, 2, and 3 be summarized? Option: A. A guide on photo gallery exploration, cloud synchronization, and DSLR photography tips., B. A sequential explanation of transferring photos, securing them, and organizing them within a smartphone., C. Three independent tutorials on iPhone photo management, Google Photos app features, and manual gallery browsing., D. A step-by-step guide on the physical, cloud, and app-based photo storage and transfer across platforms. Task 6: Multi-video attribute recognition Concept: This category involves confirming and extracting the attributes (such as texture, color, function, relationship, characteristic and shape) of a specific target across multiple videos, capturing the same target in different states. Difference: It focuses on cross-video attribute state transitions, rather than static attribute recognition within a single frame. Example Questions: Question: What color predominantly features in the attire of winning players across both sporting events? Option: "A. White.", "B. Red.", "C. Blue.", "D. Green." Task 7: Joint-video counting Concept: This category involves the precise identification and statistical analysis of the same target across multiple videos. Difference: The targets are distributed across multiple videos rather than a single video, requiring cross-video unified authentication of the targets, while eliminating limitations of single-video perspectives (such as occlusion). Targets may also be abstract, such as events. Example Questions: Question: How many monitors other than cell phones are in all the videos? Option: A. 1, B. 2, C. 3, D. 4. Task 8: Cross-video entity matching Concept: This category involves making similarity judgments of entities across multiple videos with varying spatiotemporal conditions (different space and time, same space and different times, different space and different times), such as identifying criminal suspects. Difference: It emphasizes identity association across non-overlapping fields of view, overcoming the limitations of single-view tracking. Example Questions: Question: Which brand is consistently featured in both videos, reinforcing the high-performance gaming theme?" Option: "A. ASUS.", "B. Corsair.", "C. Gigabyte.", "D. RGB." Task 9: Multi-view scene understanding Concept: This category involves integrating spatiotemporal clues from multiple perspectives to reconstruct the complete causal chain and logical relationships of event development. Difference: It constructs a global event knowledge graph, rather than understanding event fragments from a single perspective. Example Questions: Question: What is the common theme across the videos that ties together the movie date, romantic dinner, family daily, and cobblestone alley walk?" Option: "A. Solitude.", "B. Companionship.", "C. Adventure.", "D. Conflict."
Task 10: Multi-video temporal reasoning Concept: This category involves integrating multiple videos and making judgments about hidden logical relationships at specific times, such as predicting the future or sorting events. Difference: It addresses the asynchronous timestamp issue in single videos, enabling global temporal modeling. Example Questions: Question: Based on the interconnections, what is the correct chronological order of the criminal case depicted across the videos? Option: A. 1-2-3-4.", "B. 4-1-2-3.", "C. 1-2-4-3.", "D. 3-1-4-2. Task 11: Joint-video spatial navigating Concept: This category involves fusing multi-view geometric information to construct a 3D spatial semantic map, supporting cross-view path planning. Difference: Spatial registration and joint reasoning of multi-source visual data. Example Questions: Question: Where are the girl’s two wardrobes located in the room across all videos? Option: A. Head to the top-right corner to enter the kitchen and stove is located in the top-right corner of the kitchen, B. Head to the top-left corner to enter the kitchen and stove is located in the top-left corner of the kitchen, C. Head to the Task 12: Video difference caption Concept: This category involves fine-grained cross-video comparison, identifying differences across multiple videos in dimensions such as event progression and object states. Difference: It emphasizes the identification of differences in dynamic processes, rather than static image comparison. Example Questions: Question: What differentiates Video 1’s organization method from Video 2 and Video 3? Option: A. Uses a rotating tray., B. Combines different gadgets., C. Focuses on vertical space., D. Emphasizes utensil storage. Task 13: Cross-video counterfactual reasoning Concept: Based on the spatiotemporal factual foundation from multiple videos, this category constructs a causal inference chain for a virtual scenario. Difference: Causal effect estimation based on multi-source observational data, overcoming the limitations of single-video causal inference. Example Questions: Question: What element remains unchanged across the videos, regardless of the setting? Option: A. Clothing style., B. Dance type., C. Footwear focus., D. Presence of music. Task 14: Cross-video object recognition Concept: This category involves fusing multi-view object features to address challenges such as occlusion and deformation, enabling consistent identity recognition of objects across videos. Difference: It leverages multi-source information to compensate for the limitations of single-video perspectives. Example Questions: Question: Which item is commonly recognized across both Videos 1 and 2 despite differences in design aesthetic? Option: A. Bamboo plants.", "B. Distressed wooden door.", "C. Umbrella stands.", "D. Metalwork tools. Task 15: Cross-video event retrieval Concept: This category involves rapidly locating segments across multiple videos that meet specific event elements, enabling effective filtering and selection. Difference: The events are distributed across different videos, rather than being contained within a single video. Example Questions: Question: Which video shows the celebration of a team after successfully leading or winning in the match?
Option: A. Video 1.", "B. Video 2.", "C. Both of them.", "D. None of them.
@@
A.2 QA Generation

### v3

@@ 4.2.1 Overall performance
4.2.2 Closed-source vs open-source models
Our comprehensive evaluation reveals a consistent and significant performance disparity between closed-source and open-source MLLMs across all hierarchical task tiers within the CVBench framework. Quantitative results from Tables 2–4 demonstrate that closed-source models (e.g., GPT-4o, Gemini-2.0-flash) achieve an average accuracy of 67.2% on complex reasoning tasks, substantially outperforming open-source counterparts (e.g., Qwen2.5-VL, LLaVA-OneVision) at 55.8%, representing an 11.4% performance gap. This divergence is particularly pronounced in temporally demanding scenarios such as multi-video temporal reasoning (M.TR), where closed-source models attain 58.7% accuracy compared to merely 22.7% for open-source models, highlighting a critical vulnerability in current publicly available architectures when handling long-range dependencies.
The observed performance chasm stems from fundamental advantages inherent in proprietary systems. Closed-source models leverage superior scalability through larger parameter counts (e.g., GPT-4o’s estimated 1.7 trillion parameters) and more extensive multimodal pretraining datasets, enabling robust context retention across extended video sequences. Furthermore, these models incorporate advanced optimization techniques including specialized attention mechanisms such as cross-video memory buffers and spatiotemporal fusion modules that explicitly model inter-video dependencies, whereas open-source alternatives often rely on adapted single-video frameworks lacking these tailored components. Additionally, commercial systems benefit from substantial computational resources during inference, facilitating sophisticated frame sampling strategies and ensemble methods that remain impractical for most open-source deployments due to hardware constraints.
These findings underscore the necessity for open-source communities to develop specialized architectures specifically designed for cross-video reasoning tasks. The performance gap not only reflects current technological disparities but also highlights the limitations of existing open-source models in integrating information across multiple video sources. Bridging this divide will require innovations in efficient cross-modal attention mechanisms and the development of large-scale multimodal pretraining protocols accessible to broader research communities, ultimately advancing the field toward more equitable and capable multimodal understanding systems.
4.2.3 Sensitivity to input perturbations
Table 5: Ablation on input perturbations.
Run ID #1 #2 #3 Average
@@ 4.2.3 Sensitivity to input perturbations
4.2.4 Frame resolution
We investigate the impact of frame resolution using Qwen2.5-VL, considering different resolutions and varying numbers of input frames, as shown in Table 6. With 8 frames, accuracy increases slightly from 46.0% at 224×224 to 46.8% at 448×448 (+0.8%), indicating that higher resolution leads to better performance. Similar trends are observed with 16 and 32 frames. Additionally, increasing the number of input frames consistently improves performance across both resolutions.
Table 6: Impact of frame number & frame resolution.
@@ 4.2.4 Frame resolution
4.2.5 Number of video instances and duration
Fig. 4 illustrates how video duration and the number of videos affect model performance (Qwen2.5-VL). Accuracy improves with video length up to 900 seconds (48.8% for 0–300s), but declines for longer videos (45.6% for 900–1200s), likely due to increased sequence complexity and context management challenges.
Figure 4: Impact of video number and video duration.
We also highlight the difficulty of multi-video reasoning. Models that perform well on single-video tasks struggle with cross-video tasks. Shown in Fig. 4, the model achieves 52.2%, 47.9%, 44.8% accuracy with two, three, four videos, showing increased complexity of multi-video tasks.
4.2.6 Qualitative analysis of reasoning-based MLLMs
Qualitative analysis is crucial for diagnosing nuanced failures in multimodal large language models (MLLMs) beyond quantitative metrics. We examine the reasoning processes of Video-R1—a reinforcement learning-based MLLM—using the CVBench framework to uncover behavioral patterns revealing architectural limitations in cross-video reasoning. Empirical observations from Fig. 5 illustrate critical performance aspects including hallucination and cognitive depth deficits. These insights complement CVBench’s quantitative results, where Video-R1 achieves only 49.2% accuracy on complex reasoning tasks versus 91.3% human performance, highlighting significant real-world applicability gaps.
@@ 4.2.6 Qualitative analysis of reasoning-based MLLMs
The left sub-figure in Fig. 5 demonstrates logical hallucination, where Video-R1 generates semantically coherent but logically inconsistent inferences with the multi-video context. For example, in a cross-video sports event association task, the model incorrectly assigns causal relationships between unrelated events (e.g., claiming an action in one video directly influences an outcome in another) despite absent spatiotemporal evidence. This failure stems from two core limitations: (1) inadequate cross-video attention mechanisms that disrupt entity state consistency across video boundaries, and (2) textual bias during inference, prioritizing linguistic priors over visual evidence. Such hallucinations intensify in tasks requiring precise spatiotemporal alignment (e.g., cross-video object tracking).
Conversely, the right sub-figure reveals cognitive depth deficits. Video-R1 struggles with hierarchical reasoning tasks demanding integration of commonsense knowledge and multi-view context. For instance, when synthesizing cooking procedures from multiple demonstration videos, the model identifies individual steps but fails to construct temporally coherent workflows. This aligns with CVBench’s finding that MLLMs exhibit surface-level understanding bias—excelling in object recognition but faltering in high-level reasoning (e.g., causal inference). These limitations amplify in asynchronous multi-camera scenarios where entities must be tracked across divergent viewpoints and timescales.
Current MLLMs like Video-R1 optimize for single-video processing, lacking modules for cross-video context retention or inter-video entity disambiguation. This gap impedes real-world deployment in dynamic, multi-source video reasoning applications.
@@ 5 Conclusion
This paper has introduced CVBench, a comprehensive diagnostic benchmark specifically designed to address the fundamental challenge of spatiotemporal relational pattern reasoning across multiple videos in multimodal large language models. By establishing a hierarchical taxonomy encompassing object association, event association, and complex reasoning tasks, CVBench provides a systematic framework for evaluating how models integrate and analyze relational patterns across diverse spatiotemporal contexts. Our extensive empirical assessment demonstrates that even state-of-the-art MLLMs exhibit significant limitations in tasks requiring consistent entity tracking, causal inference, and cross-video pattern integration, revealing critical gaps in current architectural approaches to multi-stream understanding.
The primary contribution of this work lies in the development of a standardized diagnostic framework that enables precise identification of model weaknesses, particularly in inter-video context retention and entity disambiguation within complex pattern spaces. Through its carefully curated dataset and rigorous evaluation protocol, CVBench establishes an essential tool for the pattern recognition community to benchmark progress and guide the development of next-generation architectures capable of authentic multi-stream relational reasoning. These findings underscore the necessity of developing specialized mechanisms that extend beyond single-video processing to address the unique challenges of cross-video pattern analysis.
While CVBench represents a substantial advancement in evaluating video-based multimodal reasoning, certain limitations related to task complexity and scope indicate important directions for future research. Primary efforts should focus on creating innovative cross-video attention mechanisms and memory-augmented architectures that can effectively model long-range spatiotemporal dependencies. Additionally, expanding the benchmark to incorporate more diverse real-world scenarios, including extreme conditions and domain-specific applications, will further advance the field of video pattern recognition.
CRediT authorship contribution statement
Nannan Zhu: Writing, Visualization, Validation, Software, Methodology, Conceptualization; Yonghao Dong: Software, Methodology; Teng Wang: Writing, Investigation; Xueqian Li: Data curation, Software; Shengjun Deng: Investigation, Data curation, Software; Yijia Wang: Investigation, Data curation; Zheng Hong: Investigation, Data curation; Tiantian Geng: Data curation, Investigation; Guo Niu: Validation, Supervision, Resources, Project administration; Hanyan Huang: Investigation, Funding acquisition, Formal analysis; Xiongfei Yao: Investigation, Data curation; Shuaiwei Jiao: Investigation, Data curation.
Declaration of competing interest
The authors declare no conflict of interest.
Data availability
The benchmark dataset and evaluation code are available at: https://github.com/Hokhim2/CVBench.
Acknowledgment
This work is supported by the Guangdong Basic and Applied Basic Research Foundation (2022A1515140103), the Research Projects of Ordinary Universities in Guangdong Province (2023KTSCX133), and the Guangdong Ordinary University Key Laboratory (2021KSYS008).
Appendix A Task definitions
@@
While single-video tasks assess a model’s comprehension of individual videos, CVBench evaluates cross-video relational reasoning. This critical capability enables MLLMs to advance from controlled lab settings to robust real-world deployment. CVBench introduces 15 carefully curated evaluation dimensions that collectively measure a model’s capacity for inter-video relational reasoning, each targeting a unique but complementary aspect of cross-video understanding. An overview of these tasks, their concepts, differences, and example questions is provided in Table LABEL:long_label.
@@
Task 1: Cross-video anomaly detection
Concept: The comparative analysis of multi-source video data identifies anomalous segments deviating from normal patterns through global spatiotemporal feature and behavioral pattern comparisons.
Difference: Analyzes anomaly saliency through multi-video event richness and viewpoint coherence rather than single-video perspectives.
Example questions:
Question: Which video features individuals directly engaged in construction work rather than showcasing materials or interior design? Option: A. Video 1, B. Video 2, C. Video 3, D. Video 4
Task 2: Cross-video scene recognition Concept: Fusion of semantic features from multiple videos to locate scene segments matching target spatiotemporal attributes. Difference: Focuses on cross-video differences rather than single-video similarities. Example questions:
Question: Did both videos feature a team successfully scoring a decisive goal or point during the match’s closing moments? Option: Yes/No.
Task 3: Multi-video key-action recognition
Concept: This category involves spatiotemporal alignment of motion trajectories across videos to identify differences in action execution, supporting action quality assessment and optimization.
Difference: It emphasizes modeling the spatiotemporal consistency of actions across multiple videos, rather than classifying actions within a single video.
Example questions:
Question: What is the key action difference between the presenter in Video 1 and the presenter in Video 2? Option: A. In Video 1, holding products is key, while in Video 2, gesturing with hands dominates, B. In Video 1, gesturing is emphasized more, while in Video 2, interacting with products is dominant, C. In both, sitting is a primary action, but outdoor walking is unique to Video 2, D. Video 2 uses textual prompts for interactions, while Video 1 does not.
Task 4: Cross-video event retrieval
Concept: This category involves rapidly locating segments across multiple videos that meet specific event elements, enabling effective filtering and selection.
Difference: The events are distributed across different videos, rather than being contained within a single video.
Example questions:
Question: Which video shows the celebration of a team after successfully leading or winning in the match?
Option: A. Video 1, B. Video 2, C. Both of them, D. None of them.
Task 5: Cross-video object recognition
Concept: This category involves fusing multi-view object features to address challenges such as occlusion and deformation, enabling consistent identity recognition of objects across videos.
Difference: It leverages multi-source information to compensate for the limitations of single-video perspectives.
Example questions:
Question: Which item is commonly recognized across both Videos 1 and 2 despite differences in design aesthetic?
Option: A. Bamboo plants, B. Distressed wooden door, C. Umbrella stands, D. Metalwork tools.
Task 6: Multi-video attribute recognition
Concept: This category involves confirming and extracting the attributes (such as texture, color, function, relationship, characteristic and shape) of a specific target across multiple videos, capturing the same target in different states.
Difference: It focuses on cross-video attribute state transitions, rather than static attribute recognition within a single frame.
Example questions:
Question: What color predominantly features in the attire of winning players across both sporting events?
Option: A. White, B. Red, C. Blue, D. Green.
Task 7: Joint-video counting Concept: This category involves the precise identification and statistical analysis of the same target across multiple videos. Difference: The targets are distributed across multiple videos rather than a single video, requiring cross-video unified authentication of the targets, while eliminating limitations of single-video perspectives (such as occlusion). Targets may also be abstract, such as events. Example questions:
Question: How many monitors other than cell phones are in all the videos? Option: A. 1, B. 2, C. 3, D. 4.
Task 8: Cross-video entity matching Concept: This category involves making similarity judgments of entities across multiple videos with varying spatiotemporal conditions (different space and time, same space and different times, different space and different times), such as identifying criminal suspects.
Difference: It emphasizes identity association across non-overlapping fields of view, overcoming the limitations of single-view tracking. Example questions:
Question: Which brand is consistently featured in both videos, reinforcing the high-performance gaming theme? Option: A. ASUS, B. Corsair, C. Gigabyte, D. RGB.
Task 9: Multi-view scene understanding
Concept: This category involves integrating spatiotemporal clues from multiple perspectives to reconstruct the complete causal chain and logical relationships of event development.
Difference: It constructs a global event knowledge graph, rather than understanding event fragments from a single perspective.
Example questions:
Question: What is the common theme across the videos that ties together the movie date, romantic dinner, family daily, and cobblestone alley walk? Option: A. Solitude, B. Companionship, C. Adventure, D. Conflict.
Task 10: Multi-video temporal reasoning
Concept: This category involves integrating multiple videos and making judgments about hidden logical relationships at specific times, such as predicting the future or sorting events.
Difference: It addresses the asynchronous timestamp issue in single videos, enabling global temporal modeling.
Example questions:
Question: Based on the interconnections, what is the correct chronological order of the criminal case depicted across the videos? Option: A. 1-2-3-4, B. 4-1-2-3, C. 1-2-4-3, D. 3-1-4-2.
Task 11: Joint-video spatial navigating
Concept: This category involves fusing multi-view geometric information to construct a 3D spatial semantic map, supporting cross-view path planning.
Difference: Spatial registration and joint reasoning of multi-source visual data.
Example questions:
Question: Where are the girl’s two wardrobes located in the room across all videos? Option: A. Head to the top-right corner to enter the kitchen and stove is located in the top-right corner of the kitchen, B. Head to the top-left corner to enter the kitchen and stove is located in the top-left corner of the kitchen, C. Head to the
Task 12: Video difference caption
Concept: This category involves fine-grained cross-video comparison, identifying differences across multiple videos in dimensions such as event progression and object states.
Difference: It emphasizes the identification of differences in dynamic processes, rather than static image comparison.
Example questions:
Question: What differentiates Video 1’s organization method from Video 2 and Video 3? Option: A. Uses a rotating tray, B. Combines different gadgets, C. Focuses on vertical space, D. Emphasizes utensil storage.
Task 13: Cross-video counterfactual reasoning
Concept: Based on the spatiotemporal factual foundation from multiple videos, this category constructs a causal inference chain for a virtual scenario.
Difference: Causal effect estimation based on multi-source observational data, overcoming the limitations of single-video causal inference.
Example questions:
Question: What element remains unchanged across the videos, regardless of the setting? Option: A. Clothing style, B. Dance type, C. Footwear focus, D. Presence of music.
Task 14: Joint-video summarization
Concept: This category involves extracting semantic information from multiple videos to generate event logic and descriptions that satisfy specific query requirements.
Difference: Cross-video information fusion and semantic compression, rather than simple fragment stitching.
Example questions:
Question: How can the overall narrative of Videos 1, 2, and 3 be summarized?
Option: A. A guide on photo gallery exploration, cloud synchronization, and DSLR photography tips, B. A sequential explanation of transferring photos, securing them, and organizing them within a smartphone, C. Three independent tutorials on iPhone photo management, Google Photos app features, and manual gallery browsing, D. A step-by-step guide on the physical, cloud, and app-based photo storage and transfer across platforms.
Task 15: Cross-video procedural transfer Concept: This category focuses on cross-video dynamic knowledge transfer capabilities, this approach analyzes the knowledge in the source video (such as first aid steps or mathematical problem solving methods) and adaptively refines or supplements the execution logic of similar tasks in the target video, while overcoming the interference caused by scene differences. Difference: Associating procedural semantics across videos, as a single video can only capture isolated processes and cannot achieve knowledge generalization or adaptation. Example questions:
Question: What decor themes from Video 2 can be transferred to Video 4 to enhance the bar’s elegance? Option: A. Floral arches and greenery, B. Heart motifs on walls, C. Outdoor nature-inspired pieces, D. Grand tents and fairy lights.
Appendix B QA generation